THE J = 1 ← 0 ROTATIONAL TRANSITIONS OF ^{12}CH^{+}, ^{13}CH^{+} and ^{12}CD^{+}

\underline{T. AMANO}, Department of Chemistry and Department of Physics and Astronomy, University of Waterloo, Waterloo, ON, Canada N2L 3G1.

The CH^{+} ion is the first molecular ion identified in interstellar space. Dunham^{a} detected a couple of unidentified lines in near-UV, and later Douglas and Herzberg^{b} identified them based on their laboratory observations. The electronic spectra have been investigated extensively. On the other hand, the pure rotational transitions are less extensively studied. Cernicharo et al^{c} reported the interstellar detection of the J=2-1, 3-2, and 4-3 transitions in NGC 7027. Pearson and Drouin^{d} reported the laboratory observation of the J=1-0 line of ^{12}CH^{+} at 835078.950 MHz and, based on this frequency, predicted the frequencies for ^{13}CH^{+} and CD^{+}. The predicted ^{13}CH^{+} frequency led to identification of the interstellar line^{e}. In this talk, we present a new set of measurements of the J=1-0 lines for the normal species together with the ^{13}C and D isotopic species. The overwhelming evidences obtained in our experiments support the new identifications. An extended negative glow discharge in a gas mixture of CH_{4} (∼ 0.5 mTorr) diluted in He (∼ 60 mTorr) was used for production of CH^{+} with the discharge current of about 15 mA. Axial magnetic filed up to 160 Gauss was applied. The normal species line exhibited a surprisingly large Zeeman splitting for a ^{1}Σ molecule. The ^{13}CH^{+} line showed the spinrotation hyperfine splitting, and at higher field of 150 Gauss an unresolved lineshape was exhibited due to combined hfs and Zeeman splittings. The spin-rotation splitting in the normal species was negligibly small. The CD^{+} line showed much smaller Zeeman and spin-rotation splittings, as expected. Details of the mechanism to induce such Zeeman effect and the spin-rotation interaction will be presented. The transition frequencies for these J = 1 − 0 lines are: 835137.498(20) MHz and 453521.847(20) MHz for ^{12}CH^{+} and CD^{+}, respectively. The transition frequencies for ^{13}CH^{+} are 830216.680(50) MHz (F = 3/2 − 1/2) and 830214.961(50) MHz (F = 1/2 − 1/2). The uncertainties reflect possible errors in correcting the Zeeman shifts.

aT. Dunham, Publ. Astron. Soc. Pac., 49, 26 (1937)

bA. E. Douglas and G. Herzberg, Astrophys. J. 94, 381 (1941)

cJ. Cernicharo et al., Astrophys. J., 483, L65 (1997)

dJ. C. Pearson and B. J. Drouin, Astrophys. J., 647, L83 (2006)

eE. Falgarone et al., Astrophys. J., 634, L49 (2005)